# Spherical bispectrum expansion and quadratic estimators

### v1

@@ I Introduction
Cosmological observations are made on the celestial sphere, and appear to be broadly consistent with statistical isotropy. The CMB and large-scale structure appear to be nearly Gaussian on large scales, suggesting that the primordial fluctuations were Gaussian or only perturbatively non-Gaussian. However, since the information in a Gaussian field is limited to the power spectrum, a lot of the available information potentially lies in the non-Gaussian signal. At perturbative order, this information lies mainly in the bispectrum (three-mode correlation) or trispectrum (connected four-mode correlation). The dominant blackbody trispectrum in the CMB is due to lensing, and can equivalently be thought of as the power spectrum of a quadratic estimator for the lensing defection field responsible for the signal. The correlation between the quadratic estimator and the statistically isotropic lensing field responsible for the trispectrum has the form of a bispectrum, and this bispectrum contains the key physical signal (but is not directly observable). This paper aims to present general decomposition of bispectra observed on the celestial sphere, which can be used to isolate and test distinct physical contributions to both primordial and late-time non-Gaussianity.
@@ II.1 Parity-even
When the bispectrum is a correlator of three fields that are all parity invariant, or contains an even number of parity odd fields, we say the bispectrum is parity even. A parity-even bispectrum must have $L+\ell_{1}+\ell_{2}$ even for overall statistical parity invariance; likewise, for parity-odd, the sum must be odd. We first focus on the parity-even case, where the reduced bispectrum [^2] In quantum mechanics language, $B^{ijk}_{L\ell_{1}\ell_{2}}$ already corresponds to the reduced matrix element since it is defined using the symmetries of the Wigner-Eckart theorem. $b_{L\ell_{1}\ell_{2}}$ is usually defined as
@@
The separable $L,\mathsf{M}$-dependent factor can be written in terms of the Wigner small d-matrix (for $L+\mathsf{M}$ even, Varshalovich et al. [15, p113])
@@ IV.2 Multi-field minimum variance
for $B$ in Eqs. (19) and (28). The most relevant cases of interest being that of reconstruction from polarization ($X,Y\in(E,B))$, which combines the $EE,EB$ and $BB$ pairs, and in combination with temperature $(X,Y\in(T,E,B))$ which uses all of them. Reconstruction noise curves in this later case are shown in Fig. 5. For deep experiments like CMB-S4, $EB$ typically dominates, with low-$L$ behaviour to $O(L/\ell)^{2}$ corrections given by (dropping small $\tilde{C}_{\ell}^{BB}$ terms)
@@ IV.3 Robustness
At low lensing multipoles, masking, and unavoidable inhomogeneities in the noise maps sourced by the scan strategy typically induces very strong noise anisotropy leading to a large mean field signals [26, e.g., App. B has a detailed discussion of the Planck mean field]. Spatially-varying uncorrelated noise is equivalent to point sources, but with the noise added on the beam-convolved map, so that $I_{\ell}=b_{\ell}^{-1}$ where $b_{\ell}$ are the moments of a circularly-symmetric beam. In the upper panel of Fig. 6 we show the contributions of these responses to the lensing convergence power spectrum,
@@ Appendix B Quantized Chebyshev polynomials
Seen as a function of $\mathsf{M}$, $\mathcal{T}^{L}_{m}$ is a polynomial of degree $m$ in $\mathsf{M}$, and $\mathcal{U}^{L}_{m}$ of degree $m-1$ in $\mathsf{M}$: this follows from the three-term recursion relation [15, p.93]
@@
The weight functions in the parity-odd and parity-even cases are related by [15, p113]
@@ B.1 Other results
Eq. (5) is a special case of a more general (Regge) symmetry, which allows exchange of the ‘lensing’ quantum number $M$ for $\mathsf{M}=\ell_{2}-\ell_{1}$: if we define $\mathsf{m}=(m_{1}-m_{2})/2$, then
@@ Appendix C GMV curved-sky quadratic estimators
are quantized analogs of the cosine and sine of $\theta_{L}$ in the triangle defined by the three momenta (the result for ${}_{\pm 1}F^{+}$ follows directly from the recursion relation, and for ${}_{\pm 1}F^{-}$ from the explicit form for the Wigner functions [36]). More generally, the ratios $(-1)^{s}\>_{s}F^{+}/_{0}F^{+}$ and $(-1)^{s-1}\>_{s}F^{-}/\>_{1}F^{-}$ are quantized versions of the trigonometric factors $\cos(s\theta_{L})$ and $\sin(s\theta_{L})/\sin\theta_{L}$, to which they reduce for large momenta: this may be seen from the recursion relation, which becomes that of the Chebyshev polynomials.
@@ Appendix D Large-lens limits
Here $\alpha$ and $\beta$ should be understood as the local parameters of interests as sourced by the global field. For lensing, these are the four elements of the magnification matrix [20] that describes the local remapping $\bm{\theta}\rightarrow\bm{\theta}+\bm{d}(\theta)$
$A_{ij}-\delta_{ij}=\frac{\partial d_{j}}{\partial\theta_{i}}=\begin{pmatrix}-\kappa-\gamma_{1}&-\gamma_{2}+\omega\\ -\gamma_{2}-\omega&-\kappa+\gamma_{1}\end{pmatrix}.$ (92)

### v2

@@ I Introduction
Cosmological observations are made on the celestial sphere, and appear to be broadly consistent with statistical isotropy. The CMB and large-scale structure appear to be nearly Gaussian on large scales, suggesting that the primordial fluctuations were Gaussian or only perturbatively non-Gaussian. However, since the information in a Gaussian field is limited to the power spectrum, a lot of the available information potentially lies in the non-Gaussian signal. At perturbative order, this information lies mainly in the bispectrum (three-mode correlation) or trispectrum (connected four-mode correlation). The dominant blackbody trispectrum in the CMB is due to lensing, and can equivalently be thought of as the power spectrum of a quadratic estimator for the lensing defection field responsible for the signal. The correlation between the quadratic estimator and the statistically isotropic lensing field responsible for the trixspectrum has the form of a bispectrum, and this bispectrum contains the key physical signal (but is not directly observable). This paper aims to present general decomposition of bispectra observed on the celestial sphere, which can be used to isolate and test distinct physical contributions to both primordial and late-time non-Gaussianity.
@@ II.1 Parity-even
When the bispectrum is a correlator of three fields that are all parity invariant, or contains an even number of parity odd fields, we say the bispectrum is parity even. A parity-even bispectrum must have $L+\ell_{1}+\ell_{2}$ even for overall statistical parity invariance; likewise, for parity-odd, the sum must be odd. We first focus on the parity-even case, where the reduced bispectrum [^2] In quantum mechanics language, $B^{ijk}_{L\ell_{1}\ell_{2}}$ already corresponds to the reduced matrix element since it is defined using the symmetries of the Wigner-Eckart theorem. $b_{L\ell_{1}\ell_{2}}$ is usually defined as [13]
@@
The separable $L,\mathsf{M}$-dependent factor can be written in terms of the Wigner small d-matrix (for $L+\mathsf{M}$ even, Varshalovich et al. [16, p113])
@@ IV.2 Multi-field minimum variance
for $B$ in Eqs. (19) and (28). The most relevant cases of interest being that of reconstruction from polarization ($X,Y\in(E,B))$, which combines the $EE,EB$ and $BB$ pairs, and in combination with temperature $(X,Y\in(T,E,B))$ which uses all of them. Reconstruction noise curves in this later case are shown in Fig. 5. For deep experiments like CMB-S4, $EB$ typically dominates, with low-$L$ behaviour to $O(L/\ell)^{2}$ corrections given by (dropping small $\tilde{C}_{\ell}^{BB}$ terms, and neglecting $C_{\ell}^{TE}$ in the filtering matrix, see appendix D)
@@ IV.3 Robustness
At low lensing multipoles, masking, and unavoidable inhomogeneities in the noise maps sourced by the scan strategy typically induces very strong noise anisotropy leading to a large mean field signals [27, e.g., App. B has a detailed discussion of the Planck mean field]. Spatially-varying uncorrelated noise is equivalent to point sources, but with the noise added on the beam-convolved map, so that $I_{\ell}=b_{\ell}^{-1}$ where $b_{\ell}$ are the moments of a circularly-symmetric beam. In the upper panel of Fig. 6 we show the contributions of these responses to the lensing convergence power spectrum,
@@ Appendix B Quantized Chebyshev polynomials
Seen as a function of $\mathsf{M}$, $\mathcal{T}^{L}_{m}$ is a polynomial of degree $m$ in $\mathsf{M}$, and $\mathcal{U}^{L}_{m}$ of degree $m-1$ in $\mathsf{M}$: this follows from the three-term recursion relation [16, p.93]
@@
The weight functions in the parity-odd and parity-even cases are related by [16, p113]
@@ B.1 Other results
Eq. (5) is a special case of the general Regge symmetry [11], which allows exchange of the ‘lensing’ quantum number $M$ for $\mathsf{M}=\ell_{2}-\ell_{1}$: if we define $\mathsf{m}=(m_{1}-m_{2})/2$, then
@@
For the squeezed triangles with $L\ll\ell_{1},\ell_{2}$ we have the following approximation consistent with the Regge symmetry
$(-1)^{\ell_{2}-m_{2}}{\begin{pmatrix}L&\ell_{1}&\ell_{2}\\ M&m_{1}&m_{2}\end{pmatrix}}\approx\frac{d^{L}_{M\mathsf{M}}(\theta)}{\sqrt{2\ell+1}}\quad\text{or}\quad{\begin{pmatrix}L&\ell_{1}&\ell_{2}\\ M&m_{1}&m_{2}\end{pmatrix}}\approx(-1)^{\ell-\mathsf{m}}(-1)^{\frac{\mathsf{M}-M}{2}}\frac{d^{L}_{M\mathsf{M}}(\theta)}{\sqrt{2\ell+1}},$ (72)
where $\cos\theta\equiv 2\mathsf{m}/(2\ell+1)$. This also obeys the exact $(M,\mathsf{M})\leftrightarrow(-M,-\mathsf{M})$ symmetry of Eq. (71) and hence cannot have odd $\mathcal{O}(\mathsf{M}/\ell)$ or $\mathcal{O}(M/\ell)$ corrections, giving higher-order accuracy that the original form of the approximation that does not respect the Regge symmetry [34].
@@ Appendix C GMV curved-sky quadratic estimators
are quantized analogs of the cosine and sine of $\theta_{L}$ in the triangle defined by the three momenta (the result for ${}_{\pm 1}F^{+}$ follows directly from the recursion relation, and for ${}_{\pm 1}F^{-}$ from the explicit form for the Wigner functions [38]). More generally, the ratios $(-1)^{s}\>_{s}F^{+}/_{0}F^{+}$ and $(-1)^{s-1}\>_{s}F^{-}/\>_{1}F^{-}$ are quantized versions of the trigonometric factors $\cos(s\theta_{L})$ and $\sin(s\theta_{L})/\sin\theta_{L}$, to which they reduce for large momenta: this may be seen from the recursion relation, which becomes that of the Chebyshev polynomials. In that regime, up to $(L/\ell)^{2}$ and $(s/\ell)^{2}$ corrections (but without other constraints on $L$), we may write
$(-1)^{s}{}_{s}F^{+}\approx{}_{0}F^{+}\approx(-1)^{\frac{L+2\ell}{2}}\frac{|\Delta^{L}_{0\mathsf{M}}|}{\sqrt{2\ell+1}}\quad\text{ and }\quad(-1)^{s-1}\>{}_{s}F^{-}\approx s\>{}_{1}F^{-}\approx(-1)^{\frac{L+2\ell+1}{2}}\frac{2s}{2\ell+1}\sqrt{L(L+1)}\frac{|\Delta^{L}_{1\mathsf{M}}|}{\sqrt{2\ell+1}},$ (90)
consistent with Eq. (72).
@@ Appendix D Large-lens limits
Here $\alpha$ and $\beta$ should be understood as the local parameters of interests as sourced by the global field. For lensing, these are the four elements of the magnification matrix [21] [^10] Our full-sky definitions give a different sign to $\omega$ in the flat-limit, compared to that reference. In our conventions, an image appears to the observer rotated anti-clockwise by the angle $\omega$. that describes the local remapping $\bm{\theta}\rightarrow\bm{\theta}+\bm{d}(\theta)$
$A_{ij}-\delta_{ij}=\frac{\partial d_{j}}{\partial\theta_{i}}=\begin{pmatrix}-\kappa-\gamma_{1}&-\gamma_{2}-\omega\\ -\gamma_{2}+\omega&-\kappa+\gamma_{1}\end{pmatrix}.$ (94)